OpenStudy (australopithecus):

A detective finds a murder victim in a room with constant temperature 21◦C. At 5:00am, the body’s temperature was 34.2◦C. One hour later, it was 31.2◦C. Normal body temperature is 37◦C. Assume the body’s temperature, B(t), follows Newton’s Law of Cooling. (a) Set up and solve the differential equation for B(t). so far I have the equation B(t) = 16e^(kt) I was wondering to solve for k would I set B(1) = 31.2 ? so I would get B(1) = 31.2 = 16e^(k(1)) ln(31.2/16) = k

OpenStudy (australopithecus):

does that make sense? thus, B(t) = 16e^(ln(31.2/16)(t))

OpenStudy (australopithecus):

Newton’s Law of Cooling. states that dT/dt = k(T-T_0) where T_0 is the temperature of the room

OpenStudy (australopithecus):

and T is the temperature of the object

OpenStudy (shayaan_mustafa):

Hi Australopithecus :) How are you dear ?

OpenStudy (australopithecus):

I'm alright just confused by this question I think I'm going to look for a similar question in the text book as no one seems to be able to assist me I think I may have made a mistake.

OpenStudy (shayaan_mustafa):

Collect data first. I don't think so it is a too difficult problem. Isn't it?

OpenStudy (australopithecus):

so B(0) = (T-T_0) = 37-21 = 16 Thus the coefficient 16 is correct

OpenStudy (australopithecus):

I think my k value is reasonable as well

OpenStudy (australopithecus):

but perhaps it is wrong because it is the difference between two points

OpenStudy (shayaan_mustafa):

Tell me, what are you trying to find out first?

OpenStudy (shayaan_mustafa):

I am unable to get your concept although your equation for Newton's law of cooling is right.

OpenStudy (australopithecus):

actually I don't think my k value is reasonable

OpenStudy (australopithecus):

I'm trying to figure out the time of the murder so when the body's temperature was 34.2, I an then count back from 5:00pm to figure out the time of the murder

OpenStudy (shayaan_mustafa):

\[\Large \frac{dT}{dt}=k(T-T_0)\] \[\large \int\frac{dT}{(T-T_0)}=\int kdt\] \[\large T-T_0=Ce^{kt}-------(1)\] at t=0 i.e. T(t)=T(0) \[\large T-T_0=C\] eq(1)=> \[\large T=(T-T_0)e^{kt}+T_0\]

OpenStudy (shayaan_mustafa):

@Australopithecus kindly bump your question. So that others could help you out.

OpenStudy (australopithecus):

I don't know where you are getting the + T_0 dT/dt = k(T - T_0) dT/(T - T_0) = dtk = ln|(T - T_0)| = tk + c = T - T_0 = e^(tk + c) solve for c Set t=0 16 = e^((0)k + c) ln(16) = c

OpenStudy (shayaan_mustafa):

look at equation (1) here on the left hand side I have -T_0. So when it moves to right hand side so it will be positive.

OpenStudy (australopithecus):

the equation shouldn't be set up to = temperature, it is set up to equal time

OpenStudy (shayaan_mustafa):

No. What ever you setup equation equal to. Your first aim is to find the constant k. So it doesn't matter to which equal you setup your equation.

OpenStudy (australopithecus):

so we are assuming my k value is incorrect because I have a feeling it is

OpenStudy (shayaan_mustafa):

Yes it is incorrect. I didn't calculate. But the procedure you posted is not correct dear.

OpenStudy (australopithecus):

blah I'm stuck

OpenStudy (shayaan_mustafa):

Now what my friend?

OpenStudy (shayaan_mustafa):

It is so easy. I am sure you are confusing in variables.

OpenStudy (australopithecus):

I'm sure it is but for some reason I just can't get it

OpenStudy (australopithecus):

ok I think I might have it, I set up a separate equation to solve for, based on the idea that both share the same rate. so I have for the time when it shifts from 34.2 to 31.2 over an hour K(1) = 31.2 - 21 K(1) = 13.2e^(1)k = 10.2 k = ln(10.2/13.2) therefore B(t) = e^((ln(10.2/13.2))t)

OpenStudy (australopithecus):

does that sound more reasonable?

OpenStudy (australopithecus):

probably not

OpenStudy (australopithecus):

sorry I meant B(t) = 16e^((ln(10.2/13.2))t)

OpenStudy (australopithecus):

since I used a calculator determine the answer is 3:35pm, I know that my k value is still garbage

OpenStudy (australopithecus):

so 13.2 = 16e^(kt) 34.2 = 21 + 16e^(kt) 10.2 = 16e^(k(t+1)) I solved kt = ln(13.2/16) and k = ln(10.2/16) - ln(13.2/16)

OpenStudy (australopithecus):

I think I can then just sub k into equation 13.2 = 16e^(kt) and I get -3.60 which gives me 5-3.60 = 1:36pm which is incorrect based on what i calculated with the program I downloaded

OpenStudy (australopithecus):

I used the method listed on this site to solve the problem http://mathforum.org/library/drmath/view/62887.html